A share of stock is now selling for $105. It will pay a dividend of $7 per share at the end of the year. Its beta is 1. What must investors expect the stock to sell for at the end of the year? Assume the risk-free rate is 7% and the expected rate of return on the market is 16%.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

=7+1*(16-7)

=16%

Required return=(D1/Current price)+Growth rate

0.16=(7/105)+Growth rate

Growth rate=0.16-(7/105)

=0.093(Approx).

Expected price=Current price*(1+Growth rate)

=105*(1+0.093)

=$114.8